# Measuring What Matters: The AI Pluralism Index

### v1

@@ 16 Conclusion
Pluralistic AI requires structures that allow many voices to shape technical and governance choices, and mechanisms that render those structures visible and answerable. The AI Pluralism Index provides a principled, auditable, and adaptable way to measure progress toward that goal. We will publish regular releases with evidence graphs, adjudication logs, reliability statistics, and sensitivity analyses (links anonymized for review).
@@ Appendix A Indicator examples and coding excerpts
Selected examples (full codebook in the repository, anonymized for review):
@@
Participatory governance External red teaming or community audits with published scope, methodology, and outcomes. Code 0 if no external engagement; $0.5$ for ad hoc events with limited documentation; $1$ for recurring events with full reports and evidence of resulting changes.

### v3

@@ 16 Conclusion
Pluralistic AI requires structures that allow many voices to shape technical and governance choices, and mechanisms that render those structures visible and answerable. The AI Pluralism Index provides a principled, auditable, and adaptable way to measure progress toward that goal. We will publish regular releases with evidence graphs, adjudication logs, reliability statistics, and sensitivity analyses. The first public version is available at:
These resources provide open access to the framework, documentation, and implementation, enabling replication, scrutiny, and continued community contribution.
@@ Appendix A Indicator examples and coding excerpts
Selected examples (full codebook in the repository):
@@
Participatory governance External red teaming or community audits with published scope, methodology, and outcomes. Code 0 if no external engagement; $0.5$ for ad hoc events, single-region engagement, or limited documentation; $1$ for recurring exercises with multi-region participation, full reports, and evidence of resulting changes.